# Graph Generative Pre-trained Transformer

### v1

@@ Abstract
Graph generation is a critical task in numerous domains, including molecular design and social network analysis, due to its ability to model complex relationships and structured data. While most modern graph generative models utilize adjacency matrix representations, this work revisits an alternative approach that represents graphs as sequences of node set and edge set. We advocate for this approach due to its efficient encoding of graphs and propose a novel representation. Based on this representation, we introduce the Graph Generative Pre-trained Transformer (G2PT), an auto-regressive model that learns graph structures via next-token prediction. To further exploit G2PT’s capabilities as a general-purpose foundation model, we explore fine-tuning strategies for two downstream applications: goal-oriented generation and graph property prediction. We conduct extensive experiments across multiple datasets. Results indicate that G2PT achieves superior generative performance on both generic graph and molecule datasets. Furthermore, G2PT exhibits strong adaptability and versatility in downstream tasks from molecular design to property prediction.
@@ 1 Introduction
Graph generation has become a crucial task in various domains, from chemical discovery to social network analysis, owing to graphs’ ability to represent complex relationships and produce realistic, structured data (Du et al., 2021; Zhu et al., 2022). Recent advancements in graph generative models primarily focus on permutation-invariant methods leveraging diffusion-based approaches (Ho et al., 2020; Austin et al., 2021). For example, models like EDP-GNN (Niu et al., 2020) and GDSS (Jo et al., 2022a) represent graphs using continuous adjacency matrices. In contrast, DiGress (Vignac et al., 2022) and EDGE (Chen et al., 2023) employ discrete diffusion, treating node types and all node pairs (edges) as categorical variables. Earlier neural graph generation methods introduced permutation-dependent models, such as GraphRNN (You et al., 2018b) and DeepGMG (Li et al., 2018). hese approaches employed auto-regressive frameworks (e.g., RNNs or LSTMs (Sherstinsky, 2020)) to sequentially generate graphs. For instance, GraphRNN generates adjacency matrix entries step-by-step. DeepGMG frames graph generation as a sequence of actions (e.g., add-node, add-edge), and utilizes an agent-based model to learn the action trajectories.
The discrete diffusion-based methods, in particular, gain more attention due to it perfectly aligns with the nature of discrete structure of graphs. However, sampling from discrete diffusion models relies on the conditional-independent assumption (Gu et al., 2017): the distribution of each entry at time is independent of others, given the observation at time . This may lead to a poor approximation of the true distribution, as multiple trajectories can emerge from the current state. A similar argument is made in Campbell et al. (2022), drawing a connection to the Tau-Leaping approximation, which allows changing multiple entries at a single time step. Ideally, the most accurate sampling is to change only one dimension at one timestep, or effectively, employ a large number of denoising steps. In the graph generation problem, this is equivalent to only change the label of one node, or change one node pair in terms of addition, deletion or alteration at a time. Such sampling schema naturally degenerates into an any-order auto-regressive model (Hoogeboom et al., 2021; Campbell et al., 2022).
| Model (Rep.) | Likelihood | Illustration | #Network Calls | #Variables | Decomposition |
@@
Conditional independent | |||
| Auto-regressive () | ![]() |
@@
Due to the recent success of large language models (Achiam et al., 2023; Dubey et al., 2024), in this work, we revisit the family of auto-regressive graph generative models. We propose a new method to represent graph as sequence. Specifically, our designed sequence consists of two parts: node definition and edge definition. The node definition is first established to provide information about the node index and node type. After that, the edge definition specifies how edges are connected by using the defined node indices, as well as the edge labels. Our propose representation is sparse as it only define edges that explicitly exits, contrasting to the adjacency representation. We then utilize transformer decoder to approximate the sequence distribution via the next-token prediction loss. We name it Graph Generative Pre-trained Transformer (G2PT).
We further investigate the potential of fine-tuning G2PT to perform downstream tasks such as goal-oriented generation and graph property prediction. For goal-oriented generation, we explore the rejection sampling fine-tuning and the reinforcement learning approaches, where both methods elevate the probability mass of graphs of interest in the pre-trained model distribution. For graph property prediction, we adapt the pre-trained parameters to a target task using its supervised objective.
@@
We evaluate G2PT on two categories of tasks: general graph generation tasks, including molecule and generic graph generation, and downstream tasks requiring fine-tuning, such as goal-oriented molecular generation and molecular property prediction. For general graph generation tasks, without any engineering on the architectures, loss design, or input feature augmentation, G2PT outperforms or on par with previous state-of-the-art (SOTA) baselines over seven datasets. We also study the scaling behavior of G2PT with increasing data and model scales. Furthermore, by fine-tuning G2PT towards generating molecules with target properties, we showcase that G2PT can be easily adapted to various generative tasks that requires additional alignment. Finally, supervised fine-tuning on MoleculeNet datasets demonstrates the effectiveness of G2PT’s learned representations for prediction tasks.
@@
We propose a novel sequence-based representation that efficiently encodes graphs;
@@
Our empirical result shows that G2PT achieves strong performance across divers graph generation and prediction tasks, outperforming or matching SOTA while adapting specific tasks effectively.
2 A Review of Graph Generative Models.
To connect our proposed method to prior works, this section provides an overview of existing graph generative models, focusing on their modeling variables and likelihood definitions. We emphasize diffusion and auto-regressive models, given their demonstrated superior performance. For simplicity, and without loss of generality, we assume graphs are undirected and featureless throughout this discussion. A comparison of different frameworks is detailed in Table 1.
Denote a graph as , where is the node set and represents the edge set. Apart from , adjacency matrix is also commonly used to represent edge connections. Although the adjacency matrix is denser compared to the edge set , most existing methods prefer modeling due to its structural advantages. In the following, we discuss the likelihood definitions over these two representations, their associated model decompositions, and their strength and limitations. We assume graph size is given.
2.1 Generative Modeling of Adjacency Matrix
The likelihood of defines a joint distribution over all entries in the adjacency matrix. For auto-regressive graph models (You et al., 2018b; Liao et al., 2019), the likelihood is decomposed as follows:
where the model focuses on the strictly lower-triangular portion of the matrix. Such full-condition decomposition is universal and expressive. However, as the number of variables increases quadratically with the number of nodes , the modeling complexity escalates. Accurately approximating this distribution requires a highly expressive neural network. Moreover, generating samples demands forward passes, which can be computationally intensive.
(Discrete) diffusion models (Vignac et al., 2022; Chen et al., 2023; Qin et al., 2024), on the other hand, defines a sequence of latent variables , where . The likelihood is obtained by marginalizing the intermediate from the joint distribution
assuming entries in are independent when is given. This conditional independence assumption, while lowering the computation complexity, introduces the “multi-modality problem” that limits a model’s ability to approximate the true distribution accurately (Gu et al., 2017). The accuracy of the approximation using discrete diffusion models is further discussed by Campbell et al. (2022) under the continuous time Markov chain framework. In a word, the approximation is exact only if one uses a large denoising steps during sampling, which is similar to an auto-regressive model that changes only one entry at a time 111This statement is originally claimed by Campbell et al. (2022).. Empirically, the number of denoising steps in diffusion models often exceeds the expected number of edges in the generated graphs, as operating on the adjacency matrix needs to model both edges and non-edges.
2.2 Generative Modeling of Edge Set
While the likelihood considers all node pairs in a graph as variables, the likelihood of only considers entries that are actual edges. Specifically,
where each edge is modeled by first choosing a source node and then selecting a destination node:
This decomposition is well-suited for an auto-regressive model since it doesn’t have a fix dimension even for graphs of same size. When the training graphs are sparse (i.e., ), formulating such decomposition via auto-regressive model is computationally feasible as the number of variables is linear to .
Despite these advantages, modeling the edge set has received limited attention compared to adjacency matrix-based methods. This is largely because previous efforts (Li et al., 2018) have shown less promising results.
Remark 2.1.
The architecture for modeling adjacency matrix can leverage modern sequence models such as LSTM (Sherstinsky, 2020) or Transformers (Vaswani, 2017), as the homogeneous input and output space simplifies the learning process. This design enables the model to focus on capturing sequential relationships without needing to address varying supports or perform variable-specific transformations. In contrast, action-based frameworks used in models that operate on the edge set require more complex state transition modeling (e.g., add-node, add-edge, stop). Additionally, since nodes are incrementally added to the graph, the logits’ length for edge prediction changes dynamically. To address this, previous methods rely on a edge prediction model that uses node representations as input. However, these models are often shallow networks, limiting their expressiveness.
In this work, we show that with proper sequence design and model architecture choice, modeling edge set can achieve superior performance while being efficient.
@@ 3.1 Representing Graph as Sequence
We consider modeling a graph as a sequence of actions that first generates all nodes of a graph, then the edges among them. Denote a feature graph Here is represented as a tuple , where is the node index and is the node type. And is represented as a triple , where the first two elements define the edge connection and is the edge type. For a featureless graph, the above representation can be simplified by removing the node and edge type definitions. A graph with nodes and edges can be represented as
Here is used to denote a transition from node generation to edge generation. We illustrate it in Figure 1.
Since a graph can be encoded into different sequences by varying the node permutation and the edge generation ordering. For node orderings, we index the node using a random permutation. Based on the node indices, we obtain the edge generation orderings via the reverse of a degree-based edge-removal process shown in Alg. 1. Intuitively, the reverse of such a process first constructs an “initial seed graph” and grows it by iteratively attaching nodes to it. We also explore the effectiveness of using other canonical edge orderings such as breadth-first search (BFS) and depth-first search (DFS) (details are presented in Appendix D.1).
@@ 3.2 Learning Graph Sequences via Transformer
We utilize a transformer decoder (Vaswani, 2017) for modeling the graph sequences. Unlike language models, our defined graph sequence contains tokens from different action spaces. Here we consider using a tokenizer that unifies all types of actions into one vocabulary.
Tokenization. Let be the maximum number of nodes of a graph dataset. The unified vocabulary lookup is then defined as
@@
Training loss. We use the standard language modeling loss to minimize the negative log-likelihood
where is the parameters of the model. Since the tokens in sequence are arranged based on the defined rule, the action space for each decoding step is limited to a subset. For example, when the current input token is one of the node type, its output token can only be one of the node indices. One can impose such constraint on the output logits vector at each step to improve the modeling accuracy. In our experiment, we find that an unconstrained logits space can also yield a superior performance due to the expressiveness of transformers.
4 Fine-tuning
@@ 4.1 Goal-oriented Generation
Let be the function that estimates property of a graph . In goal-oriented generation, we are interested in obtaining a new model that generates graphs whose properties are close to more often then the pre-trained model. Such a setup has a broad application in the graph generation community such as drug discovery. In this work, we explore obtaining such distribution by fine-tuning the pre-trained model. We consider rejection sampling fine-tuning (RFT) and reinforcement learning (RL) approaches.
Rejection sampling fine-tuning. This approach fine-tunes the model using its own generated samples that satisfy the desired property . We consider the case where the property is a scalar, and an acceptance function is controlled by an distance threshold .
The algorithm for generating the fine-tuning dataset via rejection sampling is shown in Alg. 2. Note that we expect the learned pre-trained model is able to generate graphs with desired property.
When the graph of interest has a low density in the model distribution, RFT becomes inefficient as it rejects most of the samples. To address this, we further propose to self-bootstrap (SBS) the RFT model to approach the target distribution. Specifically, we first define a sequence with thresholds , where . Then we can obtained a sequence of fine-tuned models by iteratively constructing fine-tuned datasets using model trained from previous threshold. The SBS algorithm combined with RFT is shown in Alg. 3.
@@
We use the notation and interchangeable as the mapping from to is deterministic. The KL divergence prevents the target model from deviating too much from the pre-trained model.
We choose Proximal Policy Optimization (PPO) (Schulman et al., 2017) to effectively train the target policy (actor model) without sacrificing stability. We first define the token-level reward:
Here is the state of the -th step in a finite trajectory (sequence). We only assign a reward when the generation is completed. The value function of state under a model is the expectation of the undiscounted future return:
A critic model is then learned to approximate the true value function via minimizing the mean absolute error . We parameterize the critic model with a transformer that is the same architecture as the pre-trained model, except the logits head is replaced by a value head. The parameters of the critic model are also initialized from the pre-trained model.
We use the clipped surrogate objective in PPO to optimize the actor model. Moreover, to mitigate possible model degradation, we incorporate the pre-training loss following Zheng et al. (2023); Liu et al. (2024).
@@ 4.2 Property Prediction
Assume a labeled graph dataset , where each instance consists a graph along with a label . We fine-tune the pre-trained model on it to learn to predict given . After the sequence s is generated from , we extract the activation of the final token output by the last transformer block as the graph representation. To predict , we then feed into a dropout layer followed by a linear layer:
We then maximize the log-likelihood . Compared to freezing the whole transformer during training and only update parameters of the linear layer, we found that unlocking the latter half of the transformer blocks significantly enhances performance.

### v2

@@ Abstract
Graph generative models, which can produce complex structures that resemble real-world data, serve as essential tools in domains such as molecular design and network analysis. While many existing generative models rely on adjacency matrices, this work introduces a token-based approach that represents graphs as token sequences and generates them via next-token prediction, offering a more efficient encoding. Based on this methodology, we propose the Graph Generative Pre-trained Transformer (G2PT), an auto-regressive Transformer architecture that learns graph structures through this sequence-based paradigm. To extend G2PT’s capabilities as a general-purpose foundation model, we further develop fine-tuning strategies for two downstream tasks: goal-oriented generation and graph property prediction. Comprehensive experiments on multiple datasets demonstrate G2PT’s superior performance in both generic graph generation and molecular generation. Additionally, the good experiment results also show that G2PT can be effectively applied to goal-oriented molecular design and graph representation learning. The code of G2PT is released at https://github.com/tufts-ml/G2PT.
@@ 1 Introduction
Graph generation has emerged as a crucial task across diverse fields such as chemical discovery and social network analysis, thanks to its ability to model complex relationships and produce realistic, structured data (Du et al., 2021; Zhu et al., 2022).
| Model (Rep.) | Likelihood | Illustration | #Network Calls | #Variables | or |
@@
Intractable | |||
| Sequential() | ![]() |
@@
Early generation methods such as DeepGMG (Li et al., 2018) and GraphRNN (You et al., 2018b) model graphs with sequential models. These approaches employed sequential frameworks (e.g., RNNs or LSTMs (Sherstinsky, 2020)) to generate graphs sequentially. For instance, GraphRNN generates adjacency matrix entries step by step. For undirected graphs, it only needs to generate the lower triangular part of the adjacency matrix. DeepGMG frames graph generation as a sequence of actions (e.g., add-node, add-edge), and utilizes an agent-based model to learn the action trajectories.
Recent advances in graph generative models have primarily focused on permutation-invariant methods, particularly diffusion-based approaches (Ho et al., 2020; Austin et al., 2021). For example, models like EDP-GNN (Niu et al., 2020) and GDSS (Jo et al., 2022a) learn from adjacency matrices as continuous values. DiGress (Vignac et al., 2022) and EDGE (Chen et al., 2023) employ discrete diffusion, treating node types and all node pairs (edges and non-edges) as categorical variables. These models start from a random or a fixed adjacency matrix and run “denoising” steps to sample an adjacency matrix from the target graph distribution. They specify exchangeable (permutation-invariant) distributions over graphs by assigning the same probability to adjacency matrices of the same graph. However, achieving the permutation-invariant property has a price: the underlying neural network needs to be permutation-invariant as well, limiting the architecture choice to graph neural networks only. Discrete diffusion has an additional limitation: it samples matrix entries independently at each denoising step, making it challenging to learn the true distribution when the number of denoising steps is insufficient (Lezama et al., 2022; Campbell et al., 2022).
In recent years, the revolutionary success of large language models (Achiam et al., 2023; Dubey et al., 2024) shows the power of autoregressive Transformers and also inspired the application of these models in other fields such as image generation (Esser et al., 2021). In this work, we revisit the sequential approach to graph generation and introduce a novel token-based encoding scheme for representing sparse graphs as sequences. This new encoding strategy unlocks the potential of Transformer architectures for graph generation. We train autoregressive Transformers to generate graphs by predicting token sequences, resulting in our proposed method: Graph Generative Pre-trained Transformer (G2PT). While G2PT does not maintain permutation invariance, we argue that its capacity to learn accurate graph distributions from large-scale data outweighs this limitation. Table 1 gives a comparison of diffusion-based models, sequential models based on adjacency matrices, and our model based on edges.
This new approach offers an additional advantage: it allows us to leverage recent training techniques developed for Transformers in NLP. Specifically, we apply our model to goal-oriented generation tasks, such as optimizing molecular properties. To this end, we explore both rejection sampling fine-tuning and reinforcement learning, each designed to increase the probability mass assigned to desirable graphs. Additionally, our model can be used to learn graph representations when provided with the entire graph as input – a capability not feasible with diffusion-based generative models. With fine-tuning and a supervised objective, G2PT can also be adapted for graph-level tasks such as graph classification.
@@
We evaluate G2PT on a series of graph generation tasks: generic graph generation, molecule generation, and goal-oriented molecular generation. Without excessive architectural engineering or training tricks, G2PT performs better than or on par with previous state-of-the-art (SOTA) baselines over seven datasets. By fine-tuning G2PT towards generating molecules with target properties, we showcase that G2PT can be easily adapted to various generative tasks that require additional alignment. We also fine-tune G2PT for molecular property prediction on MoleculeNet datasets. The results demonstrate the effectiveness of G2PT’s learned representations for classification tasks. Finally, we analyze the G2PT model and show its performance with data augmentation.
@@
We propose a novel token-based graph representation that enables efficient graph generation;
@@
We conduct an extensive empirical study of G2PT, which achieves strong performance across diverse graph generation and prediction tasks.
2 Related work
Permutation-invariant vs sequential models. Neural graph generative models are first developed as autoregressive models. These models represent graphs as sequences and learn to generate such representations. A variety of models such as GraphRNN (You et al., 2018a), GRAN (Liao et al., 2019), GraphDF (Luo et al., 2021), and DAGG (Han et al., 2023) generate entries of the lower triangular of the adjacency matrix of the target graph, while DeepGMG (Li et al., 2018) and BiGG (Dai et al., 2020) directly generate edges through node representations learned from partially generated graphs.
In parallel, another research direction considers permutation-invariant models to predict logits of adjacency matrices and then sample the adjacency matrix in one step (Madhawa et al., 2019; Liu et al., 2019). These models define exchangeable distributions over graphs. This approach has been further advanced with diffusion-based methods (Ho et al., 2020; Austin et al., 2021), which can be divided into continuous diffusion models (Niu et al., 2020; Jo et al., 2022b) and discrete diffusion models (Vignac et al., 2022; Chen et al., 2023; Qin et al., 2024). Among these, discrete diffusion models have shown superior performance due to their alignment with the discrete nature of graphs. These models begin with a fixed or random adjacency matrix and iteratively update matrix entries over multiple steps to converge to the target adjacency matrix. Each update step, referred to as a ”denoising” step, modifies a subset of the adjacency matrix entries.
Discrete diffusion graph models typically require many denoising steps for graph sampling. This is because entries are sampled independently conditioned on the previous step for each step. Such a sampling scheme introduces the compounding decoding error (Lezama et al., 2022), leading to a poor approximation to the true distribution (Campbell et al., 2022).
Generating adjacency matrices vs edge lists. Most graph generative models, including the majority of sequential approaches (You et al., 2018b; Liao et al., 2019) and all diffusion-based methods (Jo et al., 2022b; Niu et al., 2020; Vignac et al., 2022; Chen et al., 2023; Qin et al., 2024), generate graphs via their adjacency matrix, with a few exceptions (Li et al., 2018; Dai et al., 2020) that operate directly on the edge list of the target graph.
Despite the popularity, learning adjacency matrices requires modeling all node pairs in a graph, and thus needs a computation quadratic in the number of nodes . For autoregressive models, it creates long sequences for the model to learn, making it a challenging learning problem (Hihi & Bengio, 1995). Therefore, generating adjacency matrices are computationally intensive for both autoregressive and diffusion-based models. In contrast, edge lists only define variables over actual edges. Compared to adjacency matrices, models that use edge lists usually require less computation for both learning and sampling, especially when dealing with sparse graphs.
Despite the advantage mentioned above, modeling edge lists has received limited attention because of the difficulty of modeling structurally heterogeneous discrete events (e.g., adding a node and deciding the two ends of an edge). Li et al. (2018) propose an action-based framework and model a graph with pre-defined actions such as “add-node” and “add-edge”. With a special design of the model architecture, the model is challenging to train. Additionally, the “add-edge” action cannot be treated as label prediction because the number of nodes increases with the generation procedure. Previous methods rely on node representations learned from graph neural networks, which are shown to have limited expressiveness in general (Li & Leskovec, 2022). This work takes a different approach and designs a token-based sequences that enable generic Transformer architectures.
@@ 3.1 Representing Graph as Sequence
We consider modeling a graph as a sequence that first lists all nodes and then all edges. Let denote a graph where nodes and edges are each associated with a type label. Here is represented as a tuple , where is the node index and is the node type. And is represented as a triple , where the first two elements define the edge connection and is the edge type. For a graph without node or edge labels, the above representation can be simplified by removing the node or edge labels from the sequence. A graph with nodes and edges can be represented as
Here is the special token separating node tokens from edge tokens. We illustrate it in Figure 1.
The sequence depends on a node order and an edge order. Nodes are ordered randomly. Nodes in the first part of the sequence follow the node order, and thus . The edge order is determined by reversing a degree-based edge-removal process, as described in Alg. 1. The removal process prioritizes the removal of edges connected to low-degree nodes, so its reverse constructs a compact, relatively dense core first, followed by the addition of edges with fewer connections. Edges in the second part of the sequence follow this construction order. We also explore the effectiveness of using other edge orderings such as breadth-first search (BFS) and depth-first search (DFS) (details are presented in Appendix D.1).
@@ 3.2 Learning Graph Sequences via Transformer
We utilize a Transformer decoder (Vaswani, 2017) for modeling the graph sequences. Before this, we construct a unified vocabulary containing all previously introduced tokens. A tokenizer is defined to map each token to a unique integer.
Tokenization. Let be the maximum number of nodes of a graph dataset. The unified vocabulary is then defined as
@@
Training objective. With the autoregressive Transformer, we minimize the negative log-likelihood of the sequence.
where denotes model parameters. With the rules of forming sequences from graphs, not all tokens are legal for at a prediction step . For example, when the current token is a node type, the next token can only be one of the node indices. Illegal tokens can be avoided by masking out their logits in the model output. However, our experiments show that unconstrained logits also yield superior performance thanks to the learning power of Transformers.
Relationship with graph likelihood. We show that maximizing the sequence likelihood maximizes a lower bound of the graph likelihood. Denote a training set of graphs . The log-likelihood of the graph dataset is . Let denote the data distribution, and denote the distribution of drawing sequences from a graph . We have
In the third line, we use the fact that because determines . Note that both the entropy terms are constant with respect to the model parameters .
From the lower bound, we see that an accurate approximation can tighten the bound. Although is not directly compute, providing more sequences from the same graph can reduce the condition KL. Intuitively, increasing the number of sequences increases the diversity of the training data, thus improving the model’s generalization. Combining graph datasets and random sequences, we use the objective (1) to pre-train a Transformer model as the generative model.
4 Fine tuning
@@ 4.1 Goal-oriented Generation
Let denote the function that estimates property of a graph . In goal-oriented generation, our objective is to train a new model that generates graphs with property values closer to a target value than those produced by the pre-trained model. This problem has a broad application such as drug discovery. In this work, we fine-tune the pre-trained model to improve its ability to generate graphs that better meet the specified property criteria. To this end, we explore two strategies: rejection sampling fine-tuning (RFT) and reinforcement learning (RL).
Rejection sampling fine-tuning. This approach fine-tunes the model using its generated samples that have the desired property values. Here we consider the case that the property is a scalar and specify and an acceptance function , where the distance tolerance is a hyperparameter.
We run the generative model and collect valid graphs that satisfy the acceptance function to construct the fine-tuning dataset . The algorithm is shown in Alg. 2. Note that we expect the learned pre-trained model to have the ability to generate a decent fraction of graphs with the desired property.
RFT becomes inefficient if the pre-trained model can rarely generate acceptable samples. To address this, we run the self-bootstrap (SBS) version of RFT to approach the target distribution in multiple rounds. With tolerances , we obtain a sequence of fine-tuned models by iteratively constructing fine-tuned datasets using the model trained from the previous tolerance. The SBS algorithm combined with RFT is shown in Alg. 3.
@@
Here as uniquely decides . The KL divergence prevents the target model from deviating too much from the pre-trained model.
We choose Proximal Policy Optimization (PPO) (Schulman et al., 2017) to effectively train the target model without sacrificing stability. The token-level reward is only at the last token and zero otherwise:
Here is the state of the -th step in a finite trajectory (sequence). The value function of state under a model is the expectation of the undiscounted future return:
A critic model is then learned to approximate the true value function via minimizing the mean absolute error . We parameterize the critic model using a Transformer with the same architecture as the pre-trained model, except that the logits head is replaced with a value head. The parameters of the critic model are also initialized from the pre-trained model.
We use the clipped surrogate objective in PPO to optimize the actor model. Moreover, to mitigate possible model degradation, we incorporate the pre-training loss following Zheng et al. (2023) and Liu et al. (2024).
@@ 4.2 Property Prediction
In a graph-level learning task, we often need to predict the label of a graph . With the pre-trained model, we fine-tune it to address the graph-level learning task. After the sequence s is generated from , we extract the activation of the final token from the last Transformer block as the graph representation. The rationale is that the model must have learned the information of the entire graph in order to expand it with further edges. We replace the token-predicting layer with an MLP to predict based on :
We minimize the cross-entropy loss over a graph classification dataset to fine-tune the model. Compared to freezing the whole Transformer during training and only updating parameters of the linear layer, we found that unlocking the latter half of the Transformer blocks significantly enhances the performance.